# Implementing Grassroots Logic Programs with Multiagent Transition Systems and AI

### v1

@@ 5.1 Multiagent GLP
    Cold-call $p$ to $q$: A binary transaction with participants $p\neq q\in P$ where the network output stream in $c_{p}$ has a new message msg(q,X)(q,X), $c^{\prime}_{p}$ is the result of advancing the network output stream in $c_{p}$, and $c^{\prime}_{q}$ is the result of adding $X?$ to the network input stream in $c_{q}$.
Note that Communicate may be unary or binary, depending on whether $p=q$. Both Communicate and Cold-call transfer assignments from writers to readers: Communicate operates between agents sharing a paired reader and writer, while Cold-call operates through the network streams established in each agent’s initial configuration, enabling the creation of paired variables among previously-disconnected agents.
@@ 5.2 Implementing Multiagent GLP with Deterministic Agents
  • Stage 0: At boot time, the coordinator globalizes the initial goal, creating local pairs at each agent connected by global links. Agent $p$ gets $\texttt{client1}(\mathsf{Xs_{p}})$ and spawns $\texttt{global\_send}(\mathsf{Xs_{p}}?,\_w(p,1),q)$; agent $q$ gets $\texttt{monitor}(\mathsf{Xs_{q}}?)$ and creates entry $(\mathsf{Xs_{q}},p,1)$.
  • Stage 1: After $p$ assigns $\mathsf{Xs}:=[\mathtt{add}|\mathsf{Xs1}]$, the global_send goal fires, sending the value to $q$. A new link is created for $\mathsf{Xs1}$ with its own global_send goal at $p$ and entry at $q$.
  • Stage 2: After $p$ assigns $\mathsf{Xs1}:=[\mathtt{value}(\mathsf{V?})|\mathsf{Xs2}]$. Note that $p$ sends reader $\mathsf{V?}$, so $p$ creates entry $(\mathsf{V},q)$ and $q$ spawns $\texttt{global\_send}(\mathsf{V_{q}}?,\_r(p,k),p)$—the dataflow points from $q$ to $p$.
@@ Appendix B Multiagent GLP Details
maGLP Reduce is enabled for $A\in G_{p}$ iff there exists a clause $C$ with head $H$ such that $A$ and $H$ have a writer mgu $\hat{\sigma}$. Communicate applies a readers substitution $\{X?:=T\}$ to $G_{p}$; by Definition 3.7, writer mgu assigns only writers, so $A\{X?:=T\}$ and $H$ still have the same writer mgu $\hat{\sigma}$. Cold-call adds readers to $G_{q}$ but does not modify $G_{p}$’s existing goals. In both cases, the Reduce remains enabled.
@@ C.2 Global Writers Table
The global writers table $W_{p}$ of agent $p$ is an array of global writers table entries. For entries created by Globalize at index $i$, the index $i$ is the index in the global name _r(p,i)\_r(p,i). For entries created by Localize, the entry stores the remote index explicitly.
@@
The table contains only writers that await incoming assignments. When agent $p$ globalizes a reader $Y?$, it creates an entry for writer $Y$ because $p$ will receive the assignment. When agent $q$ localizes a writer global name _w(p,i)\_w(p,i), it creates an entry for its local writer because $q$ will receive the assignment. No entries are created for outgoing links—those are handled by global_send goals.
@@ C.3 The global_send Predicate
When an agent exports both ends of a variable pair (e.g., sending $[X,X?]$ to another agent, or $X$ to one agent and $X?$ to another), the Globalize operation spawns a global_send goal for the exported writer. If a value arrives on one global link and is assigned to a local writer $X$, then $X?$ becomes known, triggering any global_send goal watching $X?$. This automatically forwards the value without requiring special forwarding logic in the Receive transaction.
@@ C.5 Globalizing and Localizing Terms
    If $Y$ is a writer: allocate the next index $i$, replace $Y$ in $T_{p}^{\uparrow}$ with _w(p,i)\_w(p,i), and spawn goal $\texttt{global\_send}(Y?,\_w(p,i),q)$ into $p$’s resolvent. No entry is created—the global_send goal handles outgoing communication.
@@
    If $Y?$ is a reader: allocate the next index $i$, create entry $(Y,q)$ at index $i$ in $W^{\prime}_{p}$, and replace $Y?$ in $T_{p}^{\uparrow}$ with _r(p,i)\_r(p,i). No goal is spawned—$p$ will receive the assignment on this link.
@@
    If _w(p,i)\_w(p,i): create fresh local pair $(Y_{q},Y_{q}?)$, allocate the next index $k$ in $W^{\prime}_{q}$, add entry $(Y_{q},p,i)$, and replace _w(p,i)\_w(p,i) with $Y_{q}?$ (the reader) in $T_{q}^{\downarrow}$. No goal is spawned—$q$ will receive the assignment on this link.
@@
    If _r(p,i)\_r(p,i): create fresh local pair $(Z_{q},Z_{q}?)$, replace _r(p,i)\_r(p,i) with $Z_{q}$ (the writer) in $T_{q}^{\downarrow}$, and spawn goal $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$ into $q$’s resolvent. No entry is created—the global_send goal handles outgoing communication.
@@
  • Writer globalized at $p$: Globalize spawns $\texttt{global\_send}(Y?,\_w(p,i),q)$ at $p$. Localize adds entry $(Y_{q},p,i)$ at $q$ and puts $Y_{q}?$ in $q$’s term. When $p$ assigns $Y$, the spawned goal fires and sends the value to $q$, where the entry routes it to $Y_{q}$, making it available via $Y_{q}?$.
  • Reader globalized at $p$: Globalize adds entry $(Y,q)$ at $p$. Localize puts $Z_{q}$ in $q$’s term and spawns $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$. When $q$ assigns $Z_{q}$, the spawned goal fires and sends the value to $p$, where the entry routes it to $Y$, making it available via $Y?$.
@@
  • Writer $X$: spawns $\texttt{global\_send}(X?,\_w(p,1),q)$, no entry
  • Reader $X?$: entry $(X,q)$ at index 2, no spawn
@@
  • For _w(p,1)\_w(p,1): pair $(Y_{q},Y_{q}?)$, entry $(Y_{q},p,1)$, term gets $Y_{q}?$
  • For _r(p,2)\_r(p,2): pair $(Z_{q},Z_{q}?)$, term gets $Z_{q}$, spawns $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$
The term at $q$ is $Y_{q}?,Z_{q}]$. When $q$ assigns $Z_{q}:=T$:
@@
    $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$ fires, sends $r(p,2):=T^{\uparrow}$ to $p$
@@
    $p$ receives, finds entry $(X,q)$ at index 2, assigns $X:=T^{\downarrow}$
@@
    $\texttt{global\_send}(X?,\_w(p,1),q)$ fires, sends $w(p,1):=T^{\uparrow}$ to $q$
@@
    $q$ receives, finds entry $(Y_{q},p,1)$, assigns $Y_{q}:=T^{\downarrow}$
@@
    $Y_{q}?$ becomes known (= $T$)
The value flows from $Z_{q}$ through $p$’s local pair to $Y_{q}?$, correctly implementing the semantics where both ends of the exported pair eventually share the same value.
@@
Case $m=(\_w(p,i):=T^{\uparrow})$ where $i>0$: The message is destined for the agent that localized _w(p,i)\_w(p,i). Agent $q$ searches its global writers table for an entry $(X_{q},p,i)$ matching the remote agent $p$ and remote index $i$. Localize $T^{\uparrow}$ by $q$ to get $T_{q}^{\downarrow}$, assign $X_{q}:=T_{q}^{\downarrow}$, apply $\{X_{q}?:=T_{q}^{\downarrow}\}$ to goals containing $X_{q}?$, reactivate suspended goals, and remove the entry from $W^{\prime}_{q}$.
Case $m=(\_r(p,i):=T^{\uparrow})$: The message is destined for agent $p$ who created this global name. Agent $p$ finds entry $(X,q)$ at index $i$ in $W_{p}$. Localize $T^{\uparrow}$ by $p$ to get $T_{p}^{\downarrow}$, assign $X:=T_{p}^{\downarrow}$, apply $\{X?:=T_{p}^{\downarrow}\}$ to goals containing $X?$, reactivate suspended goals, and remove the entry from $W^{\prime}_{p}$.
@@
An entry is added to the global writers table when a global link is established with the agent as the receiver: either when globalizing a reader (expecting an incoming assignment) or when localizing a writer global name (expecting the assignment). The entry is removed when the assignment arrives and the writer is assigned. The table tracks only those writers awaiting incoming assignments from remote agents.
@@
  • Message _w(p,i):=T\_w(p,i):=T is sent from $p$ to whoever localized _w(p,i)\_w(p,i); that agent searches for an entry with remote agent $p$ and remote index $i$
  • Message _r(p,i):=T\_r(p,i):=T is sent to $p$ (the original globalizer); $p$ has an entry at index $i$ in its global writers table
@@ C.8 Variable Correspondence
  • At agent $p$: a local variable pair $(X_{p},X_{p}?)$ where both occur in $p$’s resolvent, and a spawned goal $\texttt{global\_send}(X_{p}?,\_w(p,i),q)$
  • At agent $q$: a local variable pair $(X_{q},X_{q}?)$ where both occur in $q$’s resolvent, and an entry $(X_{q},p,i)$ in $q$’s global writers table
Similarly, a global link for a reader (callback) has:
  • At agent $p$: a local pair $(Y_{p},Y_{p}?)$, and an entry $(Y_{p},q)$ at index $i$ in $p$’s global writers table
  • At agent $q$: a local pair $(Z_{q},Z_{q}?)$, and a spawned goal $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$
@@
A global link from $p$ to $q$ as in Definition C.26 implements the maGLP shared variable pair $(X,X?)$ where writer $X$ is held by $p$ and reader $X?$ is held by $q$. The correspondence is:
  • The madGLP writer $X_{p}$ corresponds to the maGLP writer $X$ at $p$
  • The madGLP reader $X_{q}?$ corresponds to the maGLP reader $X?$ at $q$
  • The readers $X_{p}?$ and writers $X_{q}$ are implementation artifacts that enable communication
@@ C.9 Correctness Proofs
    For each writer $Y$ in $T$: globalization spawns $\texttt{global\_send}(Y?,\_w(p,i),q)$ at $p$; localization creates fresh pair $(Y_{q},Y_{q}?)$ at $q$ with entry $(Y_{q},p,i)$. These form a global link implementing shared pair $(Y,Y?)$.
@@
    For each reader $Y?$ in $T$: globalization creates entry $(Y,q)$ at index $i$ in $p$’s global writers table; localization creates fresh pair $(Z_{q},Z_{q}?)$ at $q$ and spawns $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$. These form a global link implementing shared pair $(Y,Y?)$.
@@
For (1): When $Y$ is a writer in $T$, globalization replaces $Y$ with _w(p,i)\_w(p,i) and spawns $\texttt{global\_send}(Y?,\_w(p,i),q)$. Localization at $q$ sees _w(p,i)\_w(p,i), creates fresh pair $(Y_{q},Y_{q}?)$, creates entry $(Y_{q},p,i)$, and replaces _w(p,i)\_w(p,i) with $Y_{q}?$ in $T_{q}^{\downarrow}$. When $Y$ is assigned at $p$, the value becomes available via $Y?$, the global_send goal fires and sends message $w(p,i):=T^{\uparrow}$ to $q$. Agent $q$’s Receive searches for entry $(Y_{q},p,i)$ and assigns $Y_{q}:=T^{\downarrow}$, making the value available via $Y_{q}?$. This matches maGLP semantics.
For (2): When $Y?$ is a reader in $T$, globalization replaces $Y?$ with _r(p,i)\_r(p,i) and creates entry $(Y,q)$ at index $i$ in $p$’s global writers table. Localization at $q$ creates fresh pair $(Z_{q},Z_{q}?)$, spawns $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$, and replaces _r(p,i)\_r(p,i) with $Z_{q}$ in $T_{q}^{\downarrow}$. When $q$ assigns $Z_{q}:=T$, the global_send goal fires and sends $r(p,i):=T^{\uparrow}$ to $p$. Agent $p$’s Receive finds entry $(Y,q)$ at index $i$ and assigns $Y:=T^{\downarrow}$, making the value available via $Y?$. This implements callback semantics.
@@
  • For writer $X$: spawns $\texttt{global\_send}(X?,\_w(p,1),q)$, no entry, term gets _w(p,1)\_w(p,1)
  • For reader $X?$: entry $(X,q)$ at index 2, no spawn, term gets _r(p,2)\_r(p,2)
@@
  • For _w(p,1)\_w(p,1): pair $(Y_{q},Y_{q}?)$, entry $(Y_{q},p,1)$, term gets $Y_{q}?$
  • For _r(p,2)\_r(p,2): pair $(Z_{q},Z_{q}?)$, spawns $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$, term gets $Z_{q}$
The term at $q$ is $\ldots,Y_{q}?,\ldots,Z_{q},\ldots]$ with two independent pairs. Agent $q$ does not know they are connected—that connection exists only through $p$’s local pair $(X,X?)$.
When $q$ assigns $Z_{q}:=T$:
@@
    $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$ fires, sends $r(p,2):=T^{\uparrow}$ to $p$
@@
    $p$’s Receive finds entry $(X,q)$ at index 2 and assigns $X:=T^{\downarrow}$
@@
    $\texttt{global\_send}(X?,\_w(p,1),q)$ fires, sends $w(p,1):=T^{\uparrow}$ to $q$
@@
    $q$’s Receive searches for entry $(Y_{q},p,1)$ and assigns $Y_{q}:=T^{\downarrow}$
@@
    $Y_{q}?$ becomes known at $q$
The assignment to $Z_{q}$ flows through $p$’s local pair to $Y_{q}?$. The same analysis applies when $X$ and $X?$ are sent to different agents.
@@
maGLP’s binary Communicate transaction from $p$ to $q$ is correctly implemented by madGLP’s mechanism: a global_send goal fires (via Reduce), then Send, then Receive at $q$.
@@
maGLP Communicate(p,q)(p,q) transfers assignment $\{X?:=T\}$ from $\sigma_{p}$ to $G_{q}$. In madGLP: (1) $p$’s Reduce assigns $X:=T$; (2) a subsequent Reduce fires global_send(X?,G,q)(X?,G,q), adding message $(G:=T^{\uparrow},q)$ to $M_{p}$; (3) Send moves the message to the channel; (4) Receive at $q$ localizes and applies the assignment. By Lemma C.28, $\pi$ maps the result to the maGLP Communicate result.

### v2

@@ 5.1 Multiagent GLP
    Cold-call $p$ to $q$: A binary transaction with participants $p\neq q\in P$ where the network output stream in $c_{p}$ has a new message msg(q,X)(q,X), $c^{\prime}_{p}$ is the result of advancing the network output stream in $c_{p}$, and $c^{\prime}_{q}$ is the result of adding $X$ to the network input stream in $c_{q}$.
Note that Communicate may be unary or binary, depending on whether $p=q$. Communicate transfers assignments from writers to readers between agents sharing a paired reader and writer. Cold-call transfers a term with its variables to $q$ through the network streams established in each agent’s initial configuration, enabling the creation of paired variables among previously-disconnected agents. Naturally, Cold-call would be the exception, as once agents share a paired variable they can use it to communicate indefinitely; moreover, an agent with two friends (with which it shares channels) may introduce them to each other, also eschewing the need for a Cold-call.
@@ 5.2 Implementing Multiagent GLP with Deterministic Agents
  • Stage 0: At boot time, the coordinator globalizes the initial goal, creating local pairs at each agent connected by global links. Agent $p$ gets $\texttt{client1}(\mathsf{Xs_{p}})$ and spawns $\texttt{global\_send}(\mathsf{Xs_{p}}?,\_r(p,1),q)$; agent $q$ gets $\texttt{monitor}(\mathsf{Xs_{q}}?)$ and creates entry $(\mathsf{Xs_{q}},p,1)$.
  • Stage 1: After $p$ assigns $\mathsf{Xs_{p}}:=[\mathtt{add}|\mathsf{Xs1_{p}}]$, the global_send goal at $p$ fires, sending the value to $q$. A new link is created for $\mathsf{Xs1}$ with its own global_send goal at $p$ and entry at $q$.
  • Stage 2: After $p$ assigns $\mathsf{Xs1_{p}}:=[\mathtt{value}(\mathsf{V_{p}})|\mathsf{Xs2_{p}}]$. Note that $p$ sends writer $\mathsf{V}$, so $p$ creates entry $(\mathsf{V_{p}},q)$ and $q$ spawns $\texttt{global\_send}(\mathsf{V_{q}?},\_w(p,k),p)$. When monitor assigns $\mathsf{V_{q}}$, the value flows $q{\to}p$.
@@ Appendix B Multiagent GLP Details
maGLP Reduce is enabled for $A\in G_{p}$ iff there exists a clause $C$ with head $H$ such that $A$ and $H$ have a writer mgu $\hat{\sigma}$. Communicate applies a readers substitution $\{X?:=T\}$ to $G_{p}$; by Definition 3.7, writer mgu assigns only writers, so $A\{X?:=T\}$ and $H$ still have the same writer mgu $\hat{\sigma}$. Cold-call adds the term $X$ to $G_{q}$’s network input stream but does not modify $G_{p}$’s existing goals. In both cases, the Reduce remains enabled.
@@ C.2 Global Writers Table
The global writers table $W_{p}$ of agent $p$ is an array of global writers table entries. For entries created by Globalize at index $i$, the index $i$ is the index in the global name _w(p,i)\_w(p,i). For entries created by Localize, the entry stores the remote index explicitly.
@@
The table contains only writers that await incoming assignments. When agent $p$ globalizes a writer $Y$, it creates an entry for $Y$ because $p$ will receive the assignment (the remote agent owns the writer and will send the value back). When agent $q$ localizes a reader global name _r(p,i)\_r(p,i), it creates an entry for its local writer because $q$ will receive the assignment. No entries are created for outgoing links—those are handled by global_send goals.
@@ C.3 The global_send Predicate
When an agent exports both ends of a variable pair (e.g., sending $[X,X?]$ to another agent, or $X$ to one agent and $X?$ to another), the Globalize operation spawns a global_send goal for the exported reader. If a value arrives on one global link and is assigned to a local writer $X$, then $X?$ becomes known, triggering any global_send goal watching $X?$. This automatically forwards the value without requiring special forwarding logic in the Receive transaction.
@@ C.5 Globalizing and Localizing Terms
    If $Y$ is a writer: allocate the next index $i$, create entry $(Y,q)$ at index $i$ in $W^{\prime}_{p}$, and replace $Y$ in $T_{p}^{\uparrow}$ with _w(p,i)\_w(p,i). No goal is spawned—$p$ will receive the assignment on this link.
@@
    If $Y?$ is a reader: allocate the next index $i$, replace $Y?$ in $T_{p}^{\uparrow}$ with _r(p,i)\_r(p,i), and spawn goal $\texttt{global\_send}(Y?,\_r(p,i),q)$ into $p$’s resolvent. No entry is created—the global_send goal handles outgoing communication.
@@
    If _w(p,i)\_w(p,i): create fresh local pair $(Y_{q},Y_{q}?)$, replace _w(p,i)\_w(p,i) with $Y_{q}$ (the writer) in $T_{q}^{\downarrow}$, and spawn goal $\texttt{global\_send}(Y_{q}?,\_w(p,i),p)$ into $q$’s resolvent. No entry is created—the global_send goal handles outgoing communication.
@@
    If _r(p,i)\_r(p,i): create fresh local pair $(Z_{q},Z_{q}?)$, allocate the next index $k$ in $W^{\prime}_{q}$, add entry $(Z_{q},p,i)$, and replace _r(p,i)\_r(p,i) with $Z_{q}?$ (the reader) in $T_{q}^{\downarrow}$. No goal is spawned—$q$ will receive the assignment on this link.
@@
  • Writer globalized at $p$: Globalize creates entry $(Y,q)$ at $p$. Localize spawns $\texttt{global\_send}(Y_{q}?,\_w(p,i),p)$ at $q$ and puts $Y_{q}$ (writer) in $q$’s term. When $q$ assigns $Y_{q}$, the spawned goal fires and sends the value to $p$, where the entry routes it to $Y$, making it available via $Y?$.
  • Reader globalized at $p$: Globalize spawns $\texttt{global\_send}(Y?,\_r(p,i),q)$ at $p$. Localize adds entry $(Z_{q},p,i)$ at $q$ and puts $Z_{q}?$ (reader) in $q$’s term. When $p$ assigns $Y$, the spawned goal fires and sends the value to $q$, where the entry routes it to $Z_{q}$, making it available via $Z_{q}?$.
@@
  • Writer $X$: entry $(X,q)$ at index 1, no spawn, term gets _w(p,1)\_w(p,1)
  • Reader $X?$: spawns $\texttt{global\_send}(X?,\_r(p,2),q)$, no entry, term gets _r(p,2)\_r(p,2)
@@
  • For _w(p,1)\_w(p,1): pair $(Y_{q},Y_{q}?)$, spawns $\texttt{global\_send}(Y_{q}?,\_w(p,1),p)$, term gets $Y_{q}$ (writer)
  • For _r(p,2)\_r(p,2): pair $(Z_{q},Z_{q}?)$, entry $(Z_{q},p,2)$, term gets $Z_{q}?$ (reader)
The term at $q$ is $Y_{q},Z_{q}?]$. When $q$ assigns $Y_{q}:=T$:
@@
    $\texttt{global\_send}(Y_{q}?,\_w(p,1),p)$ fires, sends $w(p,1):=T^{\uparrow}$ to $p$
@@
    $p$ receives, finds entry $(X,q)$ at index 1, assigns $X:=T^{\downarrow}$
@@
    $\texttt{global\_send}(X?,\_r(p,2),q)$ fires, sends $r(p,2):=T^{\uparrow}$ to $q$
@@
    $q$ receives, finds entry $(Z_{q},p,2)$, assigns $Z_{q}:=T^{\downarrow}$
@@
    $Z_{q}?$ becomes known (= $T$)
The value flows from $Y_{q}$ through $p$’s local pair to $Z_{q}?$, correctly implementing the semantics where both ends of the exported pair eventually share the same value.
@@
Case $m=(\_w(p,i):=T^{\uparrow})$ where $i>0$: The message is destined for agent $p$ who created this global name during Globalize. Agent $p$ finds entry $(X,q)$ at index $i$ in $W_{p}$. Localize $T^{\uparrow}$ by $p$ to get $T_{p}^{\downarrow}$, assign $X:=T_{p}^{\downarrow}$, apply $\{X?:=T_{p}^{\downarrow}\}$ to goals containing $X?$, reactivate suspended goals, and remove the entry from $W^{\prime}_{p}$.
Case $m=(\_r(p,i):=T^{\uparrow})$: The message is destined for the agent that localized _r(p,i)\_r(p,i). Agent $q$ searches its global writers table for an entry $(X_{q},p,i)$ matching the remote agent $p$ and remote index $i$. Localize $T^{\uparrow}$ by $q$ to get $T_{q}^{\downarrow}$, assign $X_{q}:=T_{q}^{\downarrow}$, apply $\{X_{q}?:=T_{q}^{\downarrow}\}$ to goals containing $X_{q}?$, reactivate suspended goals, and remove the entry from $W^{\prime}_{q}$.
@@
An entry is added to the global writers table when a global link is established with the agent as the receiver: either when globalizing a writer (expecting an incoming assignment from the remote agent that received the writer) or when localizing a reader global name (expecting an incoming assignment from the globalizer). The entry is removed when the assignment arrives and the writer is assigned. The table tracks only those writers awaiting incoming assignments from remote agents.
@@
  • Message _w(p,i):=T\_w(p,i):=T is sent to $p$ (the original globalizer who exported a writer and created the entry); $p$ has an entry at index $i$ in its global writers table
  • Message _r(p,i):=T\_r(p,i):=T is sent to whoever localized _r(p,i)\_r(p,i); that agent searches for an entry with remote agent $p$ and remote index $i$
@@ C.8 Variable Correspondence
  • At agent $p$: a local variable pair $(X_{p},X_{p}?)$ where both occur in $p$’s resolvent, and an entry $(X_{p},q)$ at index $i$ in $p$’s global writers table
  • At agent $q$: a local variable pair $(X_{q},X_{q}?)$ where both occur in $q$’s resolvent, and a spawned goal $\texttt{global\_send}(X_{q}?,\_w(p,i),p)$
Similarly, a global link for a reader has:
  • At agent $p$: a local pair $(Y_{p},Y_{p}?)$, and a spawned goal $\texttt{global\_send}(Y_{p}?,\_r(p,i),q)$
  • At agent $q$: a local pair $(Z_{q},Z_{q}?)$, and an entry $(Z_{q},p,i)$ in $q$’s global writers table
@@
A global link from $p$ to $q$ as in Definition C.26 implements the maGLP shared variable pair $(X,X?)$ where writer $X$ is held by $q$ and reader $X?$ is held by $p$. The correspondence is:
  • The madGLP writer $X_{q}$ corresponds to the maGLP writer $X$ at $q$
  • The madGLP reader $X_{p}?$ corresponds to the maGLP reader $X?$ at $p$
  • The writers $X_{p}$ and readers $X_{q}?$ are implementation artifacts that enable communication
@@ C.9 Correctness Proofs
    For each writer $Y$ in $T$: globalization creates entry $(Y,q)$ at index $i$ in $p$’s global writers table; localization creates fresh pair $(Y_{q},Y_{q}?)$ at $q$ and spawns $\texttt{global\_send}(Y_{q}?,\_w(p,i),p)$. These form a global link implementing shared pair $(Y,Y?)$.
@@
    For each reader $Y?$ in $T$: globalization spawns $\texttt{global\_send}(Y?,\_r(p,i),q)$ at $p$; localization creates fresh pair $(Z_{q},Z_{q}?)$ at $q$ with entry $(Z_{q},p,i)$. These form a global link implementing shared pair $(Y,Y?)$.
@@
For (1): When $Y$ is a writer in $T$, globalization replaces $Y$ with _w(p,i)\_w(p,i) and creates entry $(Y,q)$ at index $i$ in $p$’s global writers table. Localization at $q$ sees _w(p,i)\_w(p,i), creates fresh pair $(Y_{q},Y_{q}?)$, spawns $\texttt{global\_send}(Y_{q}?,\_w(p,i),p)$, and replaces _w(p,i)\_w(p,i) with $Y_{q}$ (writer) in $T_{q}^{\downarrow}$. When $q$ assigns $Y_{q}$, the value becomes available via $Y_{q}?$, the global_send goal fires and sends message $w(p,i):=T^{\uparrow}$ to $p$. Agent $p$’s Receive finds entry $(Y,q)$ at index $i$ and assigns $Y:=T^{\downarrow}$, making the value available via $Y?$. This matches maGLP semantics where the exported writer is assigned by the remote agent.
For (2): When $Y?$ is a reader in $T$, globalization replaces $Y?$ with _r(p,i)\_r(p,i) and spawns $\texttt{global\_send}(Y?,\_r(p,i),q)$ at $p$. Localization at $q$ creates fresh pair $(Z_{q},Z_{q}?)$, creates entry $(Z_{q},p,i)$, and replaces _r(p,i)\_r(p,i) with $Z_{q}?$ (reader) in $T_{q}^{\downarrow}$. When $p$ assigns $Y$, the value becomes available via $Y?$, the global_send goal fires and sends $r(p,i):=T^{\uparrow}$ to $q$. Agent $q$’s Receive finds entry $(Z_{q},p,i)$ and assigns $Z_{q}:=T^{\downarrow}$, making the value available via $Z_{q}?$. This implements the semantics where the exported reader receives the value from the globalizer.
@@
  • For writer $X$: entry $(X,q)$ at index 1, no spawn, term gets _w(p,1)\_w(p,1)
  • For reader $X?$: spawns $\texttt{global\_send}(X?,\_r(p,2),q)$, no entry, term gets _r(p,2)\_r(p,2)
@@
  • For _w(p,1)\_w(p,1): pair $(Y_{q},Y_{q}?)$, spawns $\texttt{global\_send}(Y_{q}?,\_w(p,1),p)$, term gets $Y_{q}$ (writer)
  • For _r(p,2)\_r(p,2): pair $(Z_{q},Z_{q}?)$, entry $(Z_{q},p,2)$, term gets $Z_{q}?$ (reader)
The term at $q$ is $\ldots,Y_{q},\ldots,Z_{q}?,\ldots]$ with two independent pairs. Agent $q$ does not know they are connected—that connection exists only through $p$’s local pair $(X,X?)$.
When $q$ assigns $Y_{q}:=T$:
@@
    $\texttt{global\_send}(Y_{q}?,\_w(p,1),p)$ fires, sends $w(p,1):=T^{\uparrow}$ to $p$
@@
    $p$’s Receive finds entry $(X,q)$ at index 1 and assigns $X:=T^{\downarrow}$
@@
    $\texttt{global\_send}(X?,\_r(p,2),q)$ fires, sends $r(p,2):=T^{\uparrow}$ to $q$
@@
    $q$’s Receive finds entry $(Z_{q},p,2)$ and assigns $Z_{q}:=T^{\downarrow}$
@@
    $Z_{q}?$ becomes known at $q$
The assignment to $Y_{q}$ flows through $p$’s local pair to $Z_{q}?$. The same analysis applies when $X$ and $X?$ are sent to different agents.
@@
maGLP’s binary Communicate transaction from $p$ to $q$ is correctly implemented by madGLP’s mechanism: a global_send goal fires (via Reduce) at the writer-owner, then Send, then Receive at the reader-owner.
@@
maGLP Communicate(p,q)(p,q) transfers assignment $\{X?:=T\}$ from $\sigma_{p}$ to $G_{q}$. In madGLP: (1) the writer-owner’s Reduce assigns the local writer; (2) a subsequent Reduce fires the global_send goal watching the paired reader, adding message $(G:=T^{\uparrow})$ to the outbox; (3) Send moves the message to the channel; (4) Receive at the reader-owner finds the entry in its global writers table, localizes, and applies the assignment. By Lemma C.28, $\pi$ maps the result to the maGLP Communicate result.